# A Unified Framework for Direction and Diffuseness Estimation Using Tight-Frame Microphone Arrays

### v1

@@ Abstract
This work presents a unified framework for estimating both sound-field direction and diffuseness using practical microphone arrays with different spatial configurations. Building on covariance-based diffuseness models, we formulate a velocity-only covariance approach that enables consistent diffuseness evaluation across heterogeneous array geometries without requiring mode whitening or spherical-harmonic decomposition. Three array types — an A-format array, a rigid-sphere array, and a newly proposed tight-frame array — are modeled and compared through both simulations and measurement-based experiments. The results show that the tight-frame configuration achieves near-isotropic directional sampling and reproduces diffuseness characteristics comparable to those of higher-order spherical arrays, while maintaining a compact physical structure. We further examine the accuracy of direction-of-arrival estimation based on acoustic intensity within the same framework. These findings connect theoretical diffuseness analysis with implementable array designs and support the development of robust, broadband methods for spatial-sound-field characterization.
@@ I Introduction
From the late 1980s to 1990s, the energy-transport equation provided a physically grounded framework for describing energy propagation in rooms. Derived from Boltzmann’s transport theory, it expresses temporal variation of acoustic energy density as a balance between flux and loss terms. Applying a Fourier-type approximation to the flux term yields the acoustic diffusion equation, which models energy distribution continuously in enclosed spaces. Picaut et al. [16] developed a numerical model based on this equation, later extended by Valeau et al. [17] to include anisotropy and boundary conditions. Hanyu [18] further incorporated wall scattering and absorption coefficients to characterize room diffuseness quantitatively. These developments form the theoretical core of modern room-diffusion analysis. Cox and D’Antonio [19] organized diffuser design and diffuse-field modeling into practical methodologies, and the ISO 17497 series [20, 21] standardized measurement and evaluation methods internationally.
@@
Throughout this history, the link between diffuseness estimation and intensity measurement has been widely recognized, as clearly demonstrated by the I/E method. The p-p probe method summarized by Fahy [34, 35] estimates local energy flow and reflection distributions. Guy and Abdou [36] visualized reflections and energy flow, while de Bree et al. [37, 38] introduced pressulre-velocity sensors. Zhang et al. [39] applied this p-u probes to near-field acoustic holography. Hoshi and Hanyu [40, 41] proposed estimating intensity using paired cardioid microphones. These studies advanced an integrated view of intensity and diffuseness, linking energetic analysis with spatial measurement.

### v2

@@ Abstract
This study extends the theoretical framework of diffuseness estimation to practical microphone arrays with different spatial configurations. Building upon the covariance-based model, we formulate a velocity-only covariance framework that enables consistent diffuseness evaluation across heterogeneous array geometries without requiring mode whitening or spherical harmonic decomposition. Three types of arrays — the A-format, rigid-sphere, and a newly proposed tight-frame array — are modeled and compared through both simulation and measurement experiments. The results demonstrate that the proposed tight-frame configuration achieves isotropic sampling and reproduces diffuseness characteristics equivalent to those of higher-order spherical arrays, while maintaining a compact physical structure. We also examine the accuracy of acoustic-intensity detection within the same framework. The findings bridge theoretical diffuseness analysis and practical implementations, contributing to the design of robust sound-field measurement systems.
@@ I Introduction
From the late 1980s to 1990s, the energy-transport equation provided a physically grounded framework for describing energy propagation in rooms. Derived from Boltzmann’s transport theory, it expresses temporal variation of acoustic energy density as a balance between flux and loss terms. Applying a Fourier-type approximation to the flux term yields the acoustic diffusion equation, which models energy distribution continuously in enclosed spaces. Picaut et al. [16] developed a numerical model based on this equation, later extended by Valeau et al. [17] to include anisotropy and boundary conditions. Hanyu [18] further incorporated wall scattering and absorption coefficients to characterize room diffuseness quantitatively. These developments form the theoretical core of modern room-diffusion analysis. Following this line of study, Sakuma [19] and Omoto [20] extended the framework toward more realistic nondiffuse conditions, introducing analytical models that link the statistical treatment of scattering with the observable decay characteristics in rectangular rooms. Cox and D’Antonio [21] organized diffuser design and diffuse-field modeling into practical methodologies, and the ISO 17497 series [22, 23] standardized measurement and evaluation methods internationally.
@@
Throughout this history, the link between diffuseness estimation and intensity measurement has been widely recognized, as clearly demonstrated by the I/E method. The p-p probe method summarized by Fahy [36, 37] estimates local energy flow and reflection distributions. Guy and Abdou [38] visualized reflections and energy flow, while de Bree et al. [39, 40] introduced pressulre-velocity sensors. Zhang et al. [41] applied this p-u probes to near-field acoustic holography.
The principle of acoustic intensity measurement using oppositely oriented cardioid microphones was first proposed by Bauer [42], forming the basis of what is now known as the C-C method. This approach was later refined and systematically established by Hoshi and Hanyu [43], also Hanyu and Hoshi [44], who developed reliable formulations for practical intensity measurements in various acoustic environments. These studies advanced an integrated view of intensity and diffuseness, linking energetic analysis with spatial measurement.
@@ ACKNOWLEDGMENTS
Author Declarations
Conflict of Interest: The author declares no conflicts of interest.
Data Availability: The data that support the findings of this study are available from the corresponding author upon reasonable request.
Ethics Approval: Not applicable.